Two capacitors, C1 = 4.00

Two capacitors, C1 = 4.00
Answer #1

a) Ceq = C1 + C2 = 4 + 11 = 15 uF

b) both have same applied voltage = 9V

c) q1 = C1.V = 4 x 9 = 36 uC

q2 = C2.V = 11*9 = 99 uC
Answer #2


a) Cnet = C1 + C2

Cnet = (C1+C2) = (11+4) = 44/15 = 15